# Accelerating Volumetric Medical Image Annotation via Short-Long Memory SAM 2

### v1

@@ 4.5.4 Experiments on 1-Volume Setting
In practical scenarios, practitioners may interested in annotating a single volume and apply the finetuned model on that volume directly to annotate new data to save time. Thus, we also investigate the performance of SLM-SAM 2 under the 1-Volume setting in comparison with SAM 2. Specifically, we randomly select five different volumes for training and report the average performance of the five models for each dataset. We follow the same training and inference settings as in the 5-Volume experiments, with the same test sets. Fig. 8 shows that SLM-SAM 2 achieves better average performance (DSC: 0.66, ASSD: 15.98) than SAM 2 (DSC: 0.55, ASSD: 25.73), demonstrating that SLM-SAM 2 also provides a more accurate and robust models under the 1-Volume setting. Furthermore, Fig. 7 presents the performance on each dataset, where SLM-SAM 2 consistently outperforms SAM 2 by a significant margin.
@@ 4.5.5 Ablation Study
In this section, we present an ablation study to investigate the impact of different memory bank configurations on model performance. Specifically, we examine how varying the selection and combination of slices within the memory bank influences the results. We denote the baseline SAM 2 configuration as $\mathcal{M}_{O}$, which consists of an original memory bank containing one annotated slice and six additional recent slices. Our SLM-SAM 2 is represented as $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$, which dynamically integrates the original SAM 2 memory bank ($\mathcal{M}_{O}$) with a recent memory bank that includes the most recent slice. Furthermore, we evaluate the performance when using only the recent memory bank, considering two configurations: one with the seven most recent slices ($\mathcal{M}_{R}=7$) and another with only the single most recent slice ($\mathcal{M}_{R}=1$). In addition, we explore the combination of the original memory bank $\mathcal{M}_{O}$ with the recent bank containing the seven most recent slices, denoted as $\mathcal{M}_{O}+(\mathcal{M}_{R}=7)$.
@@
As shown in Fig. 9, it can be seen that using only recent memory (i.e., $\mathcal{M}_{R}=7$ and $\mathcal{M}_{R}=1$) significantly outperforms $\mathcal{M}_{O}$. Specifically, Fig. 10 shows that utilizing recent memory alone achieves superior performance on target-absent slices, with $\mathcal{M}_{R}=7$ and $\mathcal{M}_{R}=1$ achieving DSC of 0.78 and 0.80, respectively, compared to 0.72 for $\mathcal{M}_{O}$. A potential explanation is that retaining annotated slices in the single memory bank (e.g., $\mathcal{M}_{O}$) may lead to less robustness at object boundaries and an increased risk of over-propagation errors.
@@
Notably, as shown in Table 2, $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$ consistently outperforms $\mathcal{M}_{O}+(\mathcal{M}_{R}=7)$, with the exception of CT-Spleen, where both configurations achieve a comparable DSC of approximate 0.96. This finding aligns with our expectations, as combining short-term memory bank ($\mathcal{M}_{R}=1$) and $\mathcal{M}_{O}$ enables the model to leverage both gradual temporal information and direct attention to the most recent slice.
@@
Overall, $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$ achieves the best performance on both target-present and target-absent slices, highlighting the advance memory mechanism design of SLM-SAM 2.
@@ 5 Conclusion
In this paper, we proposed SLM-SAM 2, a novel VOS method to accelerate volumetric medical image annotation by propagating annotations from a single slice to the remaining slices in volumetric medical data. SLM-SAM 2 introduces a dynamic short-long memory module that integrates both short-term and long-term memory, along with a learnable attention fuser. We evaluated SLM-SAM 2 on organs, bones and muscles across three public datasets. Both quantitative and qualitative results have demonstrated the superior performance of SLM-SAM 2 compared to leading automatic 2D and 3D models, as well as unsupervised and supervised VOS methods, effectively alleviating over-propagation issues.

### v2

@@ 4.5.4 Experiments on 1-Volume Setting
In practical scenarios, practitioners may be interested in annotating a single volume and applying the finetuned model on that volume directly to annotate new data to save time. Thus, we also investigated the performance of SLM-SAM 2 under the 1-Volume setting in comparison with SAM 2. Specifically, we randomly selected five different volumes for training and report the average performance of the five models for each dataset. We followed the same training and inference settings as in the 5-Volume experiments, with the same test sets. Fig. 8 shows that SLM-SAM 2 achieves better average performance (DSC: 0.68, ASSD: 14.75) than SAM 2 (DSC: 0.58, ASSD: 23.52), demonstrating that SLM-SAM 2 also provides a more accurate and robust models under the 1-Volume setting. Furthermore, Fig. 7 presents the performance on each dataset, where SLM-SAM 2 consistently outperforms SAM 2 by a significant margin on most datasets.
@@ 4.5.5 Ablation Study
In this section, we present an ablation study to investigate the impact of different memory bank configurations on model performance. Specifically, we examined how varying the selection and combination of slices within the memory bank influences the results. We denote the baseline SAM 2 configuration as $\mathcal{M}_{O}$, which consists of an original memory bank containing one annotated slice and six additional recent slices. Our SLM-SAM 2 is represented as $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$, which dynamically integrates the original SAM 2 memory bank ($\mathcal{M}_{O}$) with a recent memory bank that includes the most recent slice. Furthermore, we evaluated the performance when using only the recent memory bank, considering two configurations: one with the seven most recent slices ($\mathcal{M}_{R}=7$) and another with only the single most recent slice ($\mathcal{M}_{R}=1$). In addition, we explored the combination of the original memory bank $\mathcal{M}_{O}$ with the recent bank containing the seven most recent slices, denoted as $\mathcal{M}_{O}+(\mathcal{M}_{R}=7)$.
@@
Figure 11: Performance results (DSC) of SAM 2 and SLM-SAM 2 under different initial slice options (5-Volume). SLM-SAM 2 outperforms SAM 2 on most datasets across different slice initializations, while performance within each method remain comparable under varying initial slice settings.
As shown in Fig. 9, it can be seen that using only recent memory (i.e., $\mathcal{M}_{R}=7$ and $\mathcal{M}_{R}=1$) significantly outperforms $\mathcal{M}_{O}$. Specifically, Fig. 10 shows that utilizing recent memory alone achieves superior performance on target-absent slices, with $\mathcal{M}_{R}=7$ and $\mathcal{M}_{R}=1$ achieving DSC of 0.79 and 0.80, respectively, compared to 0.73 for $\mathcal{M}_{O}$. A potential explanation is that retaining annotated slices in the single memory bank (e.g., $\mathcal{M}_{O}$) may lead to less robustness at object boundaries and an increased risk of over-propagation errors.
Figure 12: Qualitative results of error analysis in VOS. SLM-SAM 2 exhibits greater robustness than SAM 2 in handling over-propagation with gaps and intermittent object reappearances.
@@
Notably, as shown in Table 2, $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$ consistently outperforms $\mathcal{M}_{O}+(\mathcal{M}_{R}=7)$, with the exception of CT-Spleen, where both configurations achieve a comparable DSC of approximately 0.96. This finding aligns with our expectations, as combining short-term memory bank ($\mathcal{M}_{R}=1$) and $\mathcal{M}_{O}$ enables the model to leverage both gradual temporal information and direct attention to the most recent slice.
@@
Overall, $\mathcal{M}_{O}+(\mathcal{M}_{R}=1)$ achieves the best performance on both target-present and target-absent slices, highlighting the advanced memory mechanism design of SLM-SAM 2.
In addition, we observe that finetuned SAM 2 underperforms non-finetuned SAM 2 on the MRI-Pancreas dataset. This may be attributed to overfitting, a common challenge in few-shot finetuning [60, 61].
4.6 Initial Slice Sensitivity Analysis
In this project, we primarily studied label propagation from the middle slice of the object (M). However, since the middle slice may not always be optimal, we also assessed performance sensitivity to initial slice choice by adding experiments starting propagation from (1) L: the slice with the largest object area, and (2) Q: the quarter slices of the object. For Q, we selected both the first-quarter and third-quarter slices and compute the average performance to mitigate ordering effects.
Table 3: Average performance (DSC) of SAM 2 and SLM-SAM 2 across all datasets under different initial slice selection options. Best results are in bold.
Methods Q L M
SAM 2 0.7261 0.7296 0.7042
SLM-SAM 2 0.8107 0.8231 0.8390
As shown in Table 3 and Fig. 11, SLM-SAM 2 consistently outperforms SAM 2 across all three initial slice options both on average and across most datasets, demonstrating the robustness of the dynamical short-long memory module regardless of initial points. Moreover, the performance within each method is comparable across different initial slice selections.
4.7 Propagation Error Analysis
In addition to over-propagation, we also evaluated SLM-SAM 2 under other common error scenarios in propagation, comparing its performance to SAM 2.
4.7.1 Over-Propagation with Gap
Over-propagation can occur immediately after the ground truth ends or after a gap. Fig. 5 (b)-(e) and 6 have demonstrated SLM-SAM 2 handles immediate over-propagation more effectively than SAM 2. As shown in Fig. 12, in cases where SAM 2 experiences over-propagation after a gap, SLM-SAM 2 can also address this issue.
4.7.2 Intermittent Object Presence
Another challenging scenarios in VOS tasks is intermittent object presence, where the target disappears and then reappears. In medical imaging, this occurs more frequently in ultrasound and surgical videos. We qualitatively evaluated such cases in Fig. 12, where the fetal head vanishes for several slices and then reappears. It can be observed that SAM 2 struggles to recover the segmentation, while SLM-SAM 2 can handle the interruption more robustly.
4.7.3 Premature Termination
In VOS, labels may terminate prematurely, resulting in missing predicted masks at object boundaries. To quantify this error, we computed the FNSR (false negative slice ratio), and the PFNSR (premature false negative slice ratio) as follows:
$FNSR=\frac{\text{\# of false negative slices}}{\text{\# of target-present slices}}$ (11)
[…]
Table 4: Comparison of SLM-SAM 2 and SAM 2 in terms of FNSR and PFNSR.
Methods FNSR PFNSR
SAM 2 3.205 $\%$ 2.669 $\%$
SLM-SAM 2 2.244 $\%$ 2.111 $\%$
where false negative slices are target-present slices with empty predictions. Premature false negative slices are the consecutive false negative slices at the boundaries of the target-present range (i.e., from the first false negative slice up to the first target-present slice with any predicted mask, and from the last false negative slice backward to the last target-present slice with any predicted mask). Thus, PFNSR quantifies the fraction of target-present slices missed due to premature termination.
As shown in Table. 4, SLM-SAM 2 yields fewer false negative slices than SAM 2. Moreover, SLM-SAM 2 achieves a lower PFNSR, demonstrating superior robustness to premature termination at object boundaries.
4.8 Multi-label Propagation
We evaluated the ability of SLM-SAM 2 on handling multi-label propagation under three settings: (1) I+I: models trained and tested separately on each class, (2) M+I: a single model trained on combined multi-class labels but evaluated per class, and (3) M+M: a single model trained and evaluated simultaneously on all classes by annotating slices containing every target. We conducted the experiments on the MRI-Gracilis and MRI-Sartorius datasets, where each slice contains labels of gracilis and sartorius muscles.
Figure 13: Quantitative results of multi-label propagation. SLM-SAM 2 consistently outperform SAM 2 in multi-label propagation.
Fig. 13 shows that SLM-SAM 2 outperforms SAM 2 across all three settings. Moreover, performance is comparable within each model across settings, demonstrating that a single finetuned SLM-SAM 2 can effectively propagate multiple classes.
4.9 Annotation Time & Correction Analysis
To reflect the practical value of SLM-SAM 2, we evaluated the efforts needed to correct annotations after label propagation. The assessment was performed on MRI-Kidney, CT-Femur and MRI-Sartorius, comparing SLM-SAM 2 to SAM 2 in terms of the SPV (seconds per volume) and the CSR (corrected-slice ratio), which is the number of corrected slices divided by the total number of slices. Two independent annotators, both non-physician researchers with substantial experience in medical image annotation, were blinded to which model produced each mask and corrected the labels on their assigned volumes in 3D Slicer [62]. We then compute the DSC score between the corrected masks and ground truth (DSCcorr) to ensure the corrections reach equivalent quality. The average saved effort in SPV and CSR ratio are calculated as:
$SPV_{saved}=\frac{SPV_{SAM2}-SPV_{SLM-SAM2}}{SPV_{SAM2}}$ (13)
$CSR_{saved}=\frac{CSR_{SAM2}-CSR_{SLM-SAM2}}{CSR_{SAM2}}$ (14)
Table 5: Comparison of practical annotation time and corrections needed between SAM 2 and SLM-SAM 2. SPV is seconds per volume and CSR is corrected-slice ratio.
Datasets Methods SPV CSR DSCcorr
MRI-Kidney SAM 2 342.05 23.142 $\%$ 0.9735
MRI-Kidney SLM-SAM 2 203.40 17.128 $\%$ 0.9744
CT-Femur SAM 2 515.10 42.191 $\%$ 0.9585
CT-Femur SLM-SAM 2 21.93 1.587 $\%$ 0.9691
MRI-Sartorius SAM 2 134.05 44.371 $\%$ 0.9470
MRI-Sartorius SLM-SAM 2 96.70 36.755 $\%$ 0.9588
Avg. Saved Effort 60.575 $\%$ 53.574 $\%$
According to Table 5, SLM-SAM 2 reduces correction effort by an average of $60.575\%$ in SPV and $53.574\%$ in CSR compared to SAM 2. This is reasonable as SLM-SAM 2 produces more accurate initial masks than SAM 2, and thus, requires fewer manual corrections, underscoring the practical value of SLM-SAM 2. Notably, SLM-SAM 2 achieves a much lower SPV than SAM 2 on the CT-Femur dataset. This is because SAM 2 is more likely to over-propagate masks from the femur to the hip bone (as shown in Fig. 6), which results in substantially more slices requiring correction for each volume, as reflected by its much higher CSR compared to SLM-SAM 2.
4.10 Computational Efficiency Evaluation
We assessed the computational efficiency of SAM 2 and SLM-SAM 2 at inference by measuring FPS (frames per second) and peak GPU memory usage. Table 6 shows that SLM-SAM 2 exhibits slightly lower FPS and higher peak GPU usage, which is expected given its additional memory attention module and dynamic fuser.
Table 6: Comparison of computational usage between SAM 2 and SLM-SAM 2 at inference. FPS is frames per second.
Methods FPS Peak GPU Usage (GB)
SAM 2 32.634 3.148
SLM-SAM 2 27.106 3.216
@@ 5 Conclusion
In this paper, we proposed SLM-SAM 2, a novel VOS method to accelerate volumetric medical image annotation by propagating annotations from a single slice to the remaining slices in volumetric medical data. SLM-SAM 2 introduces a dynamic short-long memory module that integrates both short-term and long-term memory, along with a learnable attention fuser. We evaluated SLM-SAM 2 on organs, bones, muscles and fetal head across four public datasets. Both quantitative and qualitative results have demonstrated the superior performance of SLM-SAM 2 compared to leading automatic 2D and 3D models, as well as state-of-the-art unsupervised and supervised VOS methods, effectively alleviating over-propagation issues and other common propagation errors. Furthermore, we showed that SLM-SAM 2 reduces manual correction time by $60.575\%$, highlighting its practical value in medical image annotation.